Using substitution
y=3+x
2x+y=8
can you show me how to do this problem

Answer by rfer(16322)

y=3+x
2x+y=8
2x+3+x=8
3x=5
x=5/3
---------
2*5/3+y=8
10/3+y=8
y=8-10/3
y=24/3-10/3
y=14/3
y=4 2/3
----------
proof
y=3+x
14/3=3+5/3
14/3=9/3+5/3